OpenStudy (anonymous):

i need help with the file i add pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

pl look at it help me I WILL AWARD MEDAL!!!!

OpenStudy (anonymous):

domain: \[-9\le x \le9\] range: \[0\le y \le 9\] it represents a function because for every dependent variable there is only one unique independent variable,